# Programmer Visual Attention During Context-Aware Code Summarization

### v1

@@ 8 Conclusion & Future Work
In conclusion, we designed an eye-tracking study to analyze how programmers read source code during context-aware source code summarization. We first studied how the gaze patterns change as participants progress through tasks. We observed a statistically significant decrease in fixation counts and regression rates as participants summarize more methods in a session. This observation meant that the participant retained and internalized information about the project code base from earlier tasks. Another possibility was that the participant was bored or exhausted. To test the latter, we rated and compared the quality of summaries and found no significant change in the quality of summaries, as one would expect from exhaustion or boredom. We also found that regardless of progress, the amount of context a participant looked at in terms of lines visited, methods visited, and fixations affects the quality of summaries. This effect can be best described to be similar to a bell curve, such that the quality of a summary increases with coverage of context until a peak, after which the quality declines as the summaries become less concise. Next, we categorize and analyze the context on that the participants looked at for the longest fixation duration. We found that participants spent the longest time looking at class methods, followed by class declarations and call graph. We observed this trend to be unaffected by project name, task order, session order, or participant. Based on our study, we propose three main lines of inquiry for future work:
@@
    Based on our findings in RQ2 and RQ3, we propose that automated code summarization models may be improved by using specific areas of the project as context. Based on our findings, we recommend future studies prioritize class methods first, class declaration second, and call graph third.
@@
    We hope our dataset of context-aware summaries aids future studies in automatic code summarization. We propose using these summaries to fine-tune LLMs and generate more context-aware summaries.

### v2

@@ 8 Conclusion & Future Work
In conclusion, we designed an eye tracking study to analyze how programmers read source code during context-aware source code summarization. We first studied how the gaze patterns change as participants progress through tasks. We observed a statistically significant decrease in fixation counts and regression rates as participants summarize more methods in a session. This observation meant that the participant retained and internalized information about the project code base from earlier tasks. Another possibility was that the participant was bored or exhausted. To test the latter, we rated and compared the quality of summaries and found no significant change in the quality of summaries, as one would expect from exhaustion or boredom.
We also found that regardless of progress, the amount of context a participant looked at in terms of lines visited, methods visited, and fixations affects the quality of summaries. This effect can be best described as similar to a bell curve, such that the quality of a summary increases with coverage of context until a peak, after which the quality declines as the summaries become less concise. The outcome of this effect starts to answer some open questions about what parts of the context people need to understand the rationale behind code, and indicates that finding specific areas of context may be more important than the quantity of context.
Next, we categorize and analyze the context on that the participants looked at for the longest fixation duration. We found that participants spent the longest time looking at class and instance methods, followed by class declarations and call graph. We observed this trend to be unaffected by project name, task order, session order, or participant.
This study also contributes to a body of work answering open questions about how quickly people may learn a new codebase, by quantifying that people need to read up to 35% fewer words after summarizing methods in each session. It also adds to an ongoing discussion in the literature about the need for context to understand code by providing information about not only how much context people need, but also how often they read it.
Based on our study, we propose three main lines of inquiry for future work:
@@
    Based on our findings in RQ2 and RQ3, we propose that automated code summarization models may be improved by using specific areas of the project as context. Based on our findings, we recommend future studies prioritize class and instance methods first, class declaration second, and call graph third. Additionally, analysis of scan paths could be used to discover which reading order results in the best summaries [27].
@@
    We hope our dataset of context-aware summaries aids future studies in automatic code summarization. We propose using the results of this study as a guide for generating code summaries that describe rationale, such as showing what code context should be included as inputs for AI-based technologies that generate code summaries or helping decide how large training data window sizes should be. Additionally, we hope this dataset will be useful as a starting point for research in other areas expanding into context-aware AI-driven approaches, such as bug localization.
Reproducibility To ensure reproducibility of the results, we release all datasets and source code, along with an online reproducibility guide with instructions outlining how we produced our results via an online appendix on Zenodo: